OpenStudy (anonymous):

Solve each linear 1st ordiary differential equation. Find explicit solutions y=f(x) if possible. If initial conditions are given, find the particular solution that satisfies them. y' = y+2xe^2x y(0)=3

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

This is a good reference page I think http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx the first step is to get the equation into the form \[\Large \frac{dy}{dt} + p(t)*y = g(t)\] tell me what you get when you do that step

OpenStudy (anonymous):

before doing that don't I have to state what P and Q is

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

P= -1 and Q=2xe^2x

jimthompson5910 (jim_thompson5910):

I think you mean g(t) instead of q(t) but yes, p(t) = -1 g(t) = 2xe^(2x)

OpenStudy (anonymous):

well in my class we are using P and Q P=-1 and Q=2x2^2x

jimthompson5910 (jim_thompson5910):

ok q works as well

jimthompson5910 (jim_thompson5910):

Now you'll need to find the integrating factor `mu(t)` \[\Huge \mu(t) = e^{{}^{\int p(t)dt}}\]

jimthompson5910 (jim_thompson5910):

In your case, you're using x instead of t

OpenStudy (anonymous):

mu(x0=e^5ydx

OpenStudy (anonymous):

mu(x)*

jimthompson5910 (jim_thompson5910):

\[\Huge \mu(x) = e^{{}^{\int p(x)dx}}\] \[\Huge \mu(x) = e^{{}^{\int -1dx}}\] \[\Huge \mu(x) = e^{-x}\]

OpenStudy (anonymous):

μ this sign is called mu right

jimthompson5910 (jim_thompson5910):

yes that's the greek letter mu

OpenStudy (anonymous):

okay what i got for e above is tht right?

jimthompson5910 (jim_thompson5910):

no the integrating factor should be e^(-x) and I show how I got that above

OpenStudy (anonymous):

oh okay so that's my N(x) yeah I see what I did wrong. also the greek letter mu could also be written as N(x) righ t

jimthompson5910 (jim_thompson5910):

idk what you mean by N(x)

OpenStudy (anonymous):

nvm do not worry abotu that.

OpenStudy (anonymous):

okay so after finding mu I have to find general solution right which I do not know how to set up

jimthompson5910 (jim_thompson5910):

\[\Large y \ ' = y + 2xe^{2x}\] \[\Large y \ ' - y = 2xe^{2x}\] \[\Large \color{red}{\mu(x)}*(y \ ' - y) = \color{red}{\mu(x)}*2xe^{2x}\] \[\Large \color{red}{e^{-x}}*(y \ ' - y) = \color{red}{e^{-x}}*2xe^{2x}\] \[\Large \frac{d}{dx}\left[e^{-x}*y\right] = e^{-x}*2xe^{2x}\] \[\Large \frac{d}{dx}\left[e^{-x}*y\right] = 2xe^{-x+2x}\] \[\Large \frac{d}{dx}\left[e^{-x}*y\right] = 2xe^{x}\] \[\Large \int\frac{d}{dx}\left[e^{-x}*y\right]dx = \int 2xe^{x}dx\] \[\Large e^{-x}*y = \int 2xe^{x}dx\] I'll let you finish up

OpenStudy (anonymous):

wow that's not even close to what I did.

OpenStudy (anonymous):

so the next step will be 2 integral xe^x dx

jimthompson5910 (jim_thompson5910):

yes, use integration by parts

OpenStudy (anonymous):

parts?

jimthompson5910 (jim_thompson5910):

integration by parts \[\Large \int u dv = u*v - \int v du\]

jimthompson5910 (jim_thompson5910):

u = x ---> du = dx dv = e^x ---> v = e^x \[\LARGE \int u dv = u*v - \int v du\] \[\LARGE \int x e^x = x*e^x - \int e^x dx\] \[\LARGE \int x e^x = x*e^x - e^x+C\]

OpenStudy (anonymous):

-x(e)(y)=2 integral xe sorry I am getting confused this problem is supper long and the last one I have to deal with for now.

OpenStudy (anonymous):

u = x ---> du = dx dv = e^x ---> v = e^x ∫udv=u∗v−∫vdu this u showed is just an example right

jimthompson5910 (jim_thompson5910):

yes that's just the basic template for integration by parts http://www.mathsisfun.com/calculus/integration-by-parts.html

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so integratign the left side will give me -e^-x y+c

OpenStudy (anonymous):

the right wil give us 2e^x (x-1)

jimthompson5910 (jim_thompson5910):

I'd have the +C on the right side I'm getting \[\Large e^{-x}*y = 2x*e^x - 2e^x + C\]

jimthompson5910 (jim_thompson5910):

2e^x (x-1) is equivalent to what I have on the right side (it's factored)

OpenStudy (anonymous):

according to your example it should be on the right only

OpenStudy (anonymous):

yeah I see it now

jimthompson5910 (jim_thompson5910):

Multiply both sides by \(\Large e^x\) to isolate y \[\Large e^{-x}*y = 2x*e^x - 2e^x + C\] \[\Large e^{x}*e^{-x}*y = e^{x}*\left(2x*e^x - 2e^x + C\right)\] \[\Large y = e^{x}*2x*e^x +e^{x}*(-2e^x) + e^{x}*C\] \[\Large y = 2x*e^{2x} - 2e^{2x} + Ce^{x}\] \[\Large y(x) = 2x*e^{2x} - 2e^{2x} + Ce^{x}\]

jimthompson5910 (jim_thompson5910):

Now use y(0) = 3 to find the value of C

OpenStudy (anonymous):

what will be my general solution here?

jimthompson5910 (jim_thompson5910):

I just wrote it

OpenStudy (anonymous):

yeah sorry I was reading it. okay so x=0 and y=3

jimthompson5910 (jim_thompson5910):

the general solution will have a C in it the particular solution will have C be replaced with some fixed number

OpenStudy (anonymous):

y(x)=2x∗e2x−2e2x+Cex plugging that in your equation 3(0)=2(0)(e^2(0)) - 2e^2(0) + Ce^0

jimthompson5910 (jim_thompson5910):

yes, then simplify and solve for C

OpenStudy (anonymous):

0=0(1)-2+c 1=c?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

y(0) = 3, so the LHS should be 3

jimthompson5910 (jim_thompson5910):

3 = -2 + C

OpenStudy (anonymous):

ph okay so c=5 what is LHS angain?

jimthompson5910 (jim_thompson5910):

LHS = left hand side

jimthompson5910 (jim_thompson5910):

yes C = 5

OpenStudy (anonymous):

alright tysm!

jimthompson5910 (jim_thompson5910):

you're welcome